# AbsoluteNet: A Deep Learning Neural Network to Classify Cerebral Hemodynamic Responses of Auditory Processing

### v1

@@ 1 Introduction
Among fNIRS-specific models, CNN-based models have outperformed traditional classifiers in many fNIRS-based brain–computer interface (BCI) applications. For instance, a 3-class finger-tapping fNIRS-based BCI study achieved 92.7% accuracy with a CNN, surpassing traditional support vector machine (SVM) and shallow neural network baselines on the same data [2]. Another network called fNIRSNET utilized a dual-branch spatial-temporal CNN for classifying auditory hemodynamic responses [3]. In this network, one branch applies temporal followed by spatial convolution, while the other does the reverse. Their outputs were fused for final prediction, with HbO2 and HbR concentration change signals serving as parallel input streams. Fully convolutional networks (FCNs) and Inception-based networks have also shown promise, using filters of varying lengths to capture multiscale neural dynamics recorded using fNIRS [5]. These models exploit the spatiaotemporal characteristics of the fNIRS signals and improve generalization across subjects.
Building on these advances, we propose AbsoluteNet, a deep CNN model designed to enhance the classification of auditory processing recorded through fNIRS. This network is built on two streams of convolutional fusion blocks—one for processing spatial-temporal information and the other for temporal-spatial aggregation of information from fNIRS data. These are followed by concatenation and fusion blocks that further integrate the extracted features using separable convolution, average pooling, and a classification head. Finally, the network incorporates customized activation functions—namely, the squared, absolute, and logarithm of the absolute functions—each designed to have symmetric shapes around the y-axis. Leveraging the mentioned design implementations, AbsoluteNet aims to improve robustness and precision of single-trial neural decoding in fNIRS-based auditory processing classification task.
@@ 3.2 Training Parameters and Optimization
Training Parameters: Cross-entropy loss and the Adam optimizer were used during training with an initial learning rate of $9\times 10^{-4}$ selected through hyperparameter optimization. The dropout rate was set to 0.3 to prevent the network from overfitting. The cross-entropy loss function $\mathcal{L}$ for a binary classification problem is defined as:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right]$ (2)
where $N$ is the number of samples, $y_{i}$ is the true label, and $\hat{y}_{i}$ is the predicted probability for the positive class.
@@
Hyperparameter Optimization: A genetic algorithm (GA) was employed to optimize the learning rate, convolutional kernel sizes, and average pooling dimensions. The GA was configured with a population size of 100, 30 generations, and a mutation rate of 0.1.
@@ 5 Discussion and Conclusion
This study faces a few limitations. First, the relatively small dataset constrains the generalizability of our findings in decoding auditory processing using fNIRS. Future work will address this by evaluating the proposed deep learning framework on larger and more diverse datasets to improve model robustness and performance across subjects. Second, the current analysis relies solely on fNIRS data, which provides limited temporal resolution. Future research should explore early fusion strategies that integrate complementary modalities—such as EEG—to leverage both the fast temporal resolution of EEG and the spatial specificity of fNIRS. This multimodal approach has the potential to enhance classification accuracy and offer a more comprehensive understanding of auditory processing dynamics. An additional limitation arises from the experimental design, which could benefit from a block design so that the standard and deviant trials would have 15 seconds of complete response followed possibly by a 15 seconds rest before the start of the next trial.

### v2

@@ 1 Introduction
Among fNIRS-specific models, CNN-based models have outperformed traditional classifiers in many fNIRS-based brain–computer interface (BCI) applications. For instance, a 3-class finger-tapping fNIRS-based BCI study achieved 92.7% accuracy with a CNN, surpassing traditional support vector machine (SVM) and shallow neural network baselines on the same data [2]. Another network called fNIRSNET utilized a dual-branch spatial-temporal CNN for classifying auditory hemodynamic responses [3]. In this network, one branch applies temporal followed by spatial convolution, while the other does the reverse. Their outputs were fused for final prediction, with HbO2 and HbR concentration change signals serving as parallel input streams. Fully convolutional networks (FCNs) and Inception-based networks have also shown promise, using filters of varying lengths to capture multiscale neural dynamics recorded using fNIRS [5]. These models exploit the spatiotemporal characteristics of the fNIRS signals and improve generalization across subjects.
Building on these advances, we propose AbsoluteNet, a deep CNN model designed to enhance the classification of auditory processing recorded through fNIRS. This network is built on two streams of convolutional fusion blocks—one for processing spatial-temporal information and the other for temporal-spatial aggregation of information from fNIRS data. These are followed by concatenation and fusion blocks that further integrate the extracted features using separable convolution, average pooling, and a classification head. Finally, the network incorporates customized activation functions—namely, the squared, absolute, and logarithm of the absolute functions—each designed to have symmetric shapes around the y-axis[13]. Leveraging the mentioned design implementations, AbsoluteNet aims to improve robustness and precision of single-trial neural decoding in fNIRS-based auditory processing classification task.
@@ 3.2 Training Parameters and Optimization
Training Parameters: Cross-entropy loss and the Adam optimizer were used during training with an initial learning rate of $9\times 10^{-4}$ selected through hyperparameter optimization. The dropout rate was set to 0.3 to prevent the network from overfitting.
@@
Hyperparameter Optimization: A genetic algorithm (GA) was employed to optimize the learning rate, convolutional kernel sizes, and average pooling dimensions. The GA was configured with a population size of 100, 30 generations, and a mutation rate of 0.1. An elitist selection strategy was used, in which the individuals with the lowest validation loss (i.e., highest fitness) were iteratively selected as parents for the next generation. Crossover was implemented by randomly recombining parent parameters, and mutation was applied by randomly perturbing individual genes within predefined bounds. The fitness was evaluated using the validation loss obtained over the first fold of a 5-fold cross-validation using the model trained from scratch for each individual.
@@ 5 Discussion and Conclusion
This study faces a few limitations. First, the relatively small dataset constrains the generalizability of our findings in decoding auditory processing using fNIRS. Future work will address this by evaluating the proposed deep learning framework on larger and more diverse datasets to improve model robustness and performance across subjects. Second, the current analysis relies solely on fNIRS data, which provides limited temporal resolution. Future research should explore early fusion strategies that integrate complementary modalities, such as EEG, to leverage both the fast temporal resolution of EEG and the spatial specificity of fNIRS. This multimodal approach has the potential to enhance classification accuracy and offer a more comprehensive understanding of auditory processing dynamics. An additional limitation arises from the experimental design, which could benefit from a block design so that the standard and deviant trials would have 15 seconds of complete response followed possibly by a 15 seconds rest before the start of the next trial.